At a used - book sale, 82 books are adult books and 140 are children's books. Of the adult books, 48 are nonfiction, while 60 of the children's books are nonfiction. If a book selected at random, find the probability of selecting the books for the following. Round the answers to the three decimal places.
Answer by Alan3354(69443)

what is 1: 2(y+3)=18
It's an equation.
-----
2: 2/3c+4=8
3: 1/2(c-2)=4
===========================
They're all equations.